OpenStudy (alyssakin):

A sprinter can go from a state of resting to running at 7 m/s in just 3 seconds. What is the sprinter's average acceleration? 6.9 m/s 21 m/s 3 m/s2 2.3 m/s2

OpenStudy (sparrow2):

7/3 approximetely 2.3

OpenStudy (anonymous):

v=u+at u=0 m/s t=3s v=7 m/s a=?

OpenStudy (anonymous):

formula Distance/Time but u don't got a distance so sorry Idk

OpenStudy (alyssakin):

Thank you guys.